OpenStudy (anonymous):

How do you factor u^2-v^2+2v-1? Please explain how, not just give the answer.

OpenStudy (jamesj):

Well, can you factor first v^2 - 2v + 1 ? This should look familiar to you.

OpenStudy (jamesj):

Remember that (a+b)^2 = a^2 + 2ab + b^2

OpenStudy (jamesj):

For the record, you will need later the equation for the difference of squares: a^2 - b^2 = (a-b)(a+b)

OpenStudy (jamesj):

And since I'm not hearing from you, I'll assume that you're stuck. But a useful special case of the first formula I gave you is when b=1 (a+b)^2 = (a+1)^2 = a^2 + 2a + 1 If you still stuck what happens when b= -1? (a+b)^2 = ...

OpenStudy (anonymous):

I've never used specific formulas before. But ok, i factored down to (v-1)^2, im just not sure what to do with the remaining u^2.. or would it be -u^2?

OpenStudy (jamesj):

Right So now we have u^2 - v^2 + 2v - 1 = u^2 - ( v^2 - 2v + 1) = u^2 - (v-1)^2 a difference of squares. So now you need to factor this ....

OpenStudy (anonymous):

Ok thanks, I got (v-1+u)(v-1+u),is that correct?

OpenStudy (jamesj):

No, not quite

OpenStudy (jamesj):

both of those terms are the same, but you're factoring now a difference of squares, and the factors look like (a-b)(a+b)

OpenStudy (anonymous):

Do I need to flip the signs back when I factor in the u^2?

OpenStudy (jamesj):

You are factoring now u^2 - (v-1)^2 Hint: write for now w = v-1. Then you need to factorize u^2 - w^2

OpenStudy (jamesj):

What are the factors of u^2 - w^2 ?

OpenStudy (anonymous):

Just give me one sec to figure it out

OpenStudy (jamesj):

Look above: I gave you the formula already in terms of a and b.

OpenStudy (anonymous):

(u-w)(u+w)

OpenStudy (jamesj):

Yes.

OpenStudy (jamesj):

Now get rid of the w in your expression so you just have something in terms of the variables of the problem: u and v.

OpenStudy (anonymous):

(u-v-1)(u+v-1)

OpenStudy (jamesj):

No, be very careful with signs.

OpenStudy (anonymous):

(u-v+1)(u+v-1)

OpenStudy (jamesj):

Correct.

OpenStudy (anonymous):

Ok thank you.